# Verbalized Algorithms

### v3

@@ 4 Verbalized Clustering
Verbalized Clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity. The problem with text-based clustering is that the distance function for natural language strings is difficult to define. All approaches to this task are embedding methods. Method, dataset, and prompt details are available in the appendix Sec. B. All results are averages of 5 random seeds.
Jaccard Baseline creates a distance matrix using Jaccard distance of token sets of strings, then applies sklearn’s MDS (Kruskal, 1964) to create an embedding. Sentence Transformer (ST) Baseline uses Qwen3-8B-Embedding directly, which was selected as the top performer in the MTEB leaderboard (Muennighoff et al., 2022). VA t-STE: This is a VA based on Student-t Stochastic Triplet Embedding (Van Der Maaten and Weinberger, 2012, t-STE), which makes triplet comparison queries to LLM. Triplet comparison takes three elements $x$, $y$ and $z$ in a metric space $\|\cdot\|$ and query if $\|x-y\|<\|x-z\|$. We ask the LLM a question such as “Is X closer to Y than is to Z? X: $x$, Y: $y$, Z: zz”. For each $x\in S$, we randomly sample $k$ queries from $S^{3}$ and their results from an LLM, then generate $D$-dim embeddings by minimizing the number of violations to the results. We show the results with hyperparameters $k\in{\left\{100,200\right\}},D=8$.
@@
We evaluated the methods across four datasets. Medarxiv (Muennighoff et al., 2022) contains the titles and the categories of medical papers on MedArxiv. The category forms a cluster. IT Field Defects consists of $78$ technical IT field defect reports submitted to our internal defect tracking system (unpublished), where some reports are manually marked as duplicates of another, forming a cluster. xkcdcolors-h/l are two datasets of 100 colors sampled from (Munroe, 2010). Each entry has a non-standard color name (e.g., baby blue, vomit) and a corresponding RGB value. The difference between h and l variants is that we defined 6 ground truth clusters with sklearn.AgglomerativeClustering using a lexicographic distance that prioritizes hue and lightness in the HSL color model, respectively. Assuming that colors tend to be more strongly associated with hues, xkcdcolors-l represents an out-of-distribution (OOD) task.
We evaluate the embeddings with a score defined in Eq. 1, which is the average rank of the members of the cluster each string belongs to, when sorted by the Euclidean distance. The score is normalized to $[0,1]$, where 1 indicates that cluster members are embedded closer. Given the input $S$ and a string $s\in S$, let $C(s)$ be the ground-truth cluster such that $s\in C(s)$. Let $i=\textsc{rank}(s,s^{\prime})\in{\left\{0,1,\ldots|S|-1\right\}}$ be an integer such that, when sorted by the pairwise Euclidean distance $d(e(s),e(s^{\prime\prime}))$ between $s$ and a string $s^{\prime\prime}$ across $s^{\prime\prime}\in S$ using embedding $e$, then $d(e(s),e(s^{\prime}))$ ranks $i$-th. The score is defined as:
score $=\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
@@
Table 5 shows the results. Jaccard fared surprisingly well in IT Defects, where the input often contains long output logs whose tokens carry significant information, but it underperformed in tasks that require semantic comprehension. Sentence Transformer achieves strong performance in the standard clustering tasks such as Medarxiv and IT Defects. However, it fell short in xkcdcolors-h/l that requires information about the distance criteria. The score 0.53 in the xkcdcolors-l is close to 0.5, i.e., its distance ranking is only marginally better than a dice roll. VA t-STE outperformed traditional baselines with $k=200$. We highlight its strong generalization in OOD tasks where other approaches failed, due to its strong algorithmic bias combined with flexible LLM queries.
@@ 5 Related Work
Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023). Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary). Formalization approach to planning is also called a parse-and-solve model (Collins et al., 2022) as opposed to a LLM-as-Planner model that naively generates a plan with an LLM. Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs. To our knowledge, there is no work that simulates traditional graph search algorithms (e.g., A* (Hart et al., 1968), GBFS (Bonet and Geffner, 2001)) in a VA manner, i.e., by replacing all of its atomic operations (e.g., successor generation, heap-based priority queues, hash computation for duplicate detection) with LLM queries, although some work replace some atomic operations, especially successor generations (Katz et al., 2024).
A number of recent papers in the Algorithm with Predictions (Mitzenmacher and Vassilvitskii, 2022, ALPS) community achieved a significant success in augmenting algorithms by data-driven oracles. The particular strength of this line of work is the rigorous theoretical analysis of the error rate. VAs is a form of ALPS specialized in LLM-based oracles, thus combining ideas from ALPS community with VAs is a promising direction for future work.
Combining data-driven heuristics with existing algorithms in the classical AI tasks is also a common practice in the recent literature (e.g., classical planning Núñez-Molina et al. (2024), boolean SAT Duan et al. (2020), logistics tasks Xin et al. (2021)).
6 Conclusion, Discussion and Future Work
@@
While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including submodular maximization and pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc. The potential of this paradigm is large, and we invite various ideas for unique verbalizations and new applications. For example, can Floyd’s tortoise-and-hare cycle detection algorithm for linked lists be verbalized to perform a tautology detection? Garbage collection algorithms to detect unimportant sentences in a paper draft?
@@ A.2 Kendall-Tau Score
The Kendall Tau correlation coefficient Kendall (1945) is a non-parameteric statistic that measures how much order is preserved between two sets of rankings. Kendall-Tau score has several ways to compute ($\tau_{a},\tau_{b},\tau_{c}$) depending on the existence of ties or the nature of the observations. Here we describe $\tau_{b}$ we used in our paper.
@@ B.1 Jaccard distance
B.2 t-STE
Van Der Maaten and Weinberger (2012) proposed a heavy-tailed kernel to measure local similarities between data points. In particular, we opt to use a Student-$t$ kernel with $\alpha$ degrees of freedom by defining:
$p_{ij\ell}=\frac{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}+\left(1+\frac{\lVert x_{i}-x_{\ell}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}$
We refer to the resulting technique as t-Distributed STE (t-STE). Our implementation sets $\alpha=D-1$, where $D$ is the number of dimensions in the resulting embedding.
The formulation of the triplet probabilities is motivated by the success of unsupervised dimensionality reduction technique that also employ heavy-tailed similarity kernels (Maaten and Hinton, 2008).
B.3 LLM Query Prompts
@@
IT Defect dataset:
criteria = f"You will be given three defect reports X, Y, and Z, sent by the customers of our GPFS filesystem storage system. They may contain stack traces of various programs (python, kernel drivers, ...). Similar repetitive lines are omitted as [...]. Considering the cause of the defects (i.e. ignoring timestamp/IP/MAC addresses), are X and Z more similar than Y and Z are? "
@@
B.4 IT Defect Reports Dataset
Let $D$ denote the set of all defects. A subset $D^{\prime}\subseteq D$ was manually labeled as duplicates of several other defects in $D$. We selected 78 data points so that the final dataset contains 25 such clusters, each of size at least 2.
Example Datapoint (redacted/simplified)
The following is an example of a redacted/simplified defect record that appear in the dataset.
  "id": 347852,
  "description":
  "System login:
    EMS - X.XX.XXX.XXX root/cluster
    S6K w/ Falcon attached - lothal-qa6-[1 && 2] Code levels:
    S6k - 6230_B19
    FCM4 - 4_3_3
    Falcon - 5342 --> 5343
   Snap location:
    glogin:/u/DUMPS/R.347852/
   What I was doing:
    I was upgrading Falcon enclosure from 5342 to 5343.
   What I expected:
    ’XXXXfirmware’ will start and the Falcon will go through the upgrade process.
   What happened:
    ’mmchrimware’ completed but there are errors printed.
    The Falcon does appear to have the updated firmware.
    However, there are issues when trying to get a print
    from the Left module when running ’essenclosures’.
   Analysis:
    sg91:13:/dev/sg91::naa.500093d006585000/10910::[1.0.91.0]:/dev/sg91:XXX:5149-091:5342
    [10:48:31.258608008] XXXXXXFirmwareVer2 checking if enclosure needs to be updated
    [10:48:31.260567945] XXXXXXFirmwareVer2 firmware file found
    [10:48:31.262950549] XXXXXXFirmwareVer2 -> ...
    [10:48:31.350737778] XXXXXXFirmwareVer2 ...
    [10:48:31.362309567] XXXXXXFirmwareVer2 lothal-qa6-1-hs: [E] Error 1 updating firmware ...
    sg_ses failed: Transport error
  "duplicate_of": [],

### v4

@@ 5 Verbalized Clustering
Topic clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity. The challenge with text-based clustering is defining a metric space and a distance function. While traditional approach toward this task performs $k$-Means or other standard clustering algorithms on the distance matrix over the vector embeddings of the input strings, Verbalized Clustering instead leverage triplet comparison queries to LLM. Triplet comparison takes three elements $x$, $y$ and $z$ in a metric space $\|\cdot\|$ and query if $\|x-z\|<\|y-z\|$, i.e., we ask the LLM a question such as “Is Z similar to X than is to Y? X: $x$, Y: $y$, Z: zz”.
We propose Verbalized Clustering (Alg. 1), which combines $O(n\log n)$ approximate Delaunay graph construction (Haghiri et al., 2017), a triplet-based algorithm which was originally developed for $k$-Approximate Nearest Neighbor Search, and $O(n\log^{2}n)$ Clauset-Newman-Moore greedy modularity maximization (Clauset et al., 2004), an algorithm that clusters the set of nodes on a graph based on how densely connected the nodes are. The latter does not involve LLM calls, thus $O(n\log n)$ with a large constant factor for LLM calls dominates the runtime rather than $O(n\log^{2}n)$.
1:function Cluster(strings $S$, int $k$, int $n$):
2: Initialize a graph $G=(V,E)\leftarrow(S,\emptyset)$
3:for $0<i<k$ do
4:  Roots $x,y\leftarrow$ randomly chosen from $S$
5:  $E\leftarrow E\cup{\left\{(x,y)\right\}}$, $S^{\prime}\leftarrow S\setminus{\left\{x,y\right\}}$
6:  if $S^{\prime}\not$$\emptyset$ then
7:   $\textsc{Recurse}(x,{\left\{z\in S^{\prime}\mid\|x-z\|<\|y-z\|\right\}})$
8:   $\textsc{Recurse}(y,{\left\{z\in S^{\prime}\mid\|x-z\|>\|y-z\|\right\}})$
9:return greedy_modularity_communities($G,n$)
10:function Recurse(parent $p$, strings $S$):
[…]
13:if $S^{\prime}\not$$\emptyset$ then
14:  $\textsc{Recurse}(x,{\left\{z\in S^{\prime}\mid\|x-z\|<\|y-z\|\right\}})$
15:  $\textsc{Recurse}(y,{\left\{z\in S^{\prime}\mid\|x-z\|>\|y-z\|\right\}})$
Algorithm 1 Verbalized Clustering with strings $S$, int $k$, number of output clusters $n$
Table 6: Adjusted Rand Index across methods and benchmarks. The top results are highlighted in bold.
$k$ Medarxiv xkcdcolors-h xkcdcolors-l Qwen3-8B 20 0.16 0.42 0.10 50 0.14 0.39 0.06 Qwen3-14B 20 0.14 0.51 0.10 50 0.18 0.42 0.12 Qwen3-32B 20 0.17 0.46 0.11 50 0.20 0.46 0.16 granite-embedding-125m-english 0.13 0.18 0.00 Qwen3-Embedding-8B 0.11 0.43 0.01 all-MiniLM-L6-v2 0.11 0.19 0.01
We compare VA clustering with Sentence Transformer (ST) Baseline, which uses Qwen3-8B-Embedding directly, which was selected as the top performer in the MTEB leaderboard (Muennighoff et al., 2022). The embedding vectors are then processed by K-Means to obtain the cluster labels.
@@
We evaluated the methods across three datasets. Medarxiv (Muennighoff et al., 2022) contains the titles and the categories of medical papers on MedArxiv. The category forms a cluster. xkcdcolors-h/l are two datasets of 100 colors we created from (Munroe, 2010). Each entry has a non-standard color name (e.g., baby blue, vomit) and a corresponding RGB value. The difference between h and l variants is that we defined 6 ground truth clusters with sklearn.AgglomerativeClustering using a lexicographic distance that prioritizes hue and lightness in the ground-truth HSL color values, respectively. Assuming that colors tend to be more strongly associated with hues, xkcdcolors-l represents an out-of-distribution (OOD) task.
@@
We evaluate the accuracy using the Adjusted Rand Index (Hubert & Arabie, 1985). Table 6 shows the results. Sentence Transformer achieves strong performance in the standard clustering tasks such as Medarxiv. However, it fell short in xkcdcolors-h/l that requires information about the distance criteria. The scores in the xkcdcolors-l are close to 0, i.e., its output clusters is almost equivalent to a dice roll. In all datasets, Verbalized Clustering outperformed traditional baselines. We highlight its strong generalization in OOD tasks where other approaches failed, due to its strong algorithmic bias combined with flexible LLM queries.
6 Verbalized Submodular Maximization
Retrieval Augmented Generation (RAG) is a mainstream approach for reducing LLM hallucinations by retrieving documents that are relevant to the query from a large database and appending such documents to the prompt so that LLMs can draw answers from the documents.
While existing work tend to simply rank the documents via vector embedding similarity search, this approach fails when the task requires combining information in multiple documents to draw conclusions, because all the top-ranked documents may be similar to each other and lack crucial information. To address this limitation, existing work formulated multi-hop Q&A tasks (Yang et al., 2018) and observed that there are several types in each ground-truth retrieval (e.g., bridging, comparison, etc), differing in how they chain multiple information together. Existing work proposed various heuristic methods that additionally require fine-tuning (Park et al., 2025), based on the observations.
On the other hand, traditional embedding-based approaches to multi-hop reasoning rely on a classical notion of submodularity: Given a search space $C$ and a solution $S\subseteq C$, they assume the evaluation function $f(S)$ has diminishing returns, i.e., $\forall x\in C,S\subseteq S^{\prime}\subseteq C;f(S^{\prime}\cup{\left\{x\right\}})-f(S^{\prime})\leq f(S\cup{\left\{x\right\}})-f(S)$. This approach, however, only works in a setting that has a well-defined evaluation function $f$ that has submodularity.
We propose Verbalized Greedy Submodular Maximization (VGSM), a verbalized algorithm that follows the classical greedy algorithm that guarantees 11/e1-1/e optimality. While the notion of submodularity requires a concrete evaluation function $f$, VGSM specifies a implicit score function by giving a prompt shown in Fig. 1 to the LLM. Although LLMs are not able to produce the absolute score value $f(S\cup{\left\{x\right\}})$ when it selects a new element $x\in D\setminus S$, this is not an issue in greedy submodular maximization because it only requires a relative, contrastive measure that encodes the selection policy / preference, which originates from LLMs’ Plackett-Luce reward model.
[…]
Figure 1: The submodularity-inducing query for VGSM. Note that $X$ contains all non-selected documents $D\setminus S$.
We evaluated the approach using HotpotQA (Yang et al., 2018) (fullwiki, hard) and CUAD (Hendrycks et al., 2021) datasets. Each dataset contains a set of 3-tuples $(Q,A,D)$ where $D$ is a set of documents, $Q$ is the question and $A$ is the ground-truth answer. We assume a generative model $p(A|D,Q)=\sum_{S}p_{\phi}(A|S,Q)p_{\psi}(S|Q,D)$, where $S$ is the documents selected by the retrieval model $\psi$, and $p_{\phi}(A|S,Q)$ represents the likelihood of the answer by an answering model $\phi$. Given $\psi$ and $\phi$, we evaluate the performance of $\psi$ by $\log p_{\phi}(A|S,Q)=\sum_{t}\log p_{\phi}(A_{t}|A_{1:t-1},S,Q)$, i.e., the sum of log likelihoods over each token in the answer string $A$. This log likelihood is obtained by feeding the LLM a prompt concatenating $Q,S,A$ in order.
We compared VGSM with seven baselines. TF-IDF and BM25 are the traditional bag-of-words baselines that provide a similarity metric $f(x,y)$. It sorts documents $D$ by $f(Q,d)$ for $d\in D$, and selects $k$ most similar documents. Embedding baseline uses state-of-the-art BERT/Transformer embedding models from SentenceTransformers library. Let $e(x)$ be the embedding vector of a text string $x$. It sorts $D$ by the cosine distance between $e(Q)$ and $e(x)$ for $x\in D$, and selects the top-$k$ instances (closest to $Q$). MMR-$f$ baselines applies Maximum Marginal Relevance (MMR) rule (Carbonell & Goldstein, 1998) to the similarity metric $f$. (When $f$ is a distance, use the additive inverse.) It converts $f$ into a non-monotone submodular function $f_{MMR}(x,S)=\lambda f(x,Q)-(1-\lambda)\max_{y\in S}f(x,y)$, where $S$ is a set of already selected documents, then apply the greedy submodular maximization algorithm. We evaluate this baseline to assess the impact of submodularity on non-LLM-based scoring functions. However, since MMR is non-monotone, it loses the 11/e1-1/e optimality guarantee. Finally, Verbalized Sorting uses the Symmetrized Powersort discussed in Sec. 4 to sort $D$ and retrieve top-$k$ elements. We evaluate this in order to see the impact of using LLMs for multi-hop retrieval task without imposing submodularity.
Using each method, we retrieved $k=3,6,12$ documents and evaluated $\textsc{logp}=\log p_{\phi}(A|S,Q)$. We varied $\phi$ and $\psi$ across Qwen3 models of various sizes ($\phi\in{\left\{\text{0.6B},\text{32B}\right\}}$, $\psi\in{\left\{\text{0.6B},\text{4B},\text{32B}\right\}}$). For a faster evaluation, VGSM and Verbalized Sorting pre-selects $2k$ or $4k$ documents using the embedding baseline before running the respective algorithm.
Table 7 shows logp of various methods. Traditional TF-IDF and BM25 baselines consistently performed worse than the embedding model, as expected. MMR did not help TF-IDF, BM25 or the embedding models, which could be attributed to the lack of monotone submodularity in the MMR scheme.
For every embedding model $e$, combination with VGSM ($|D|\rightarrow{\left\{2k,4k\right\}}\rightarrow k$ documents) consistently achieved higher logp than simply selecting $k$ documents with $e$, demonstrating the effectiveness of VGSM. Verbalized Sorting performed worse than VGSM, which demonstrates the impact of the implicit submodular objective encoded by the LLMs.
On the performance of VGSM, we observed: (1) Regardless of the method, allowing a larger $k$ tends to improve logp because it increases the chance of successful retrieval. (2) Regardless of the method, using a larger answer model $\phi$ improved logp, but it did not affect the relative performance among the retrievers $\psi$. (3) When $k=12$ and pre-selecting $4k=48$ documents, VSGM with $\psi=$ Qwen3-0.6B performed worse than with $2k=24$ or with smaller $k$ (highlighted in red). This is likely because selecting from 48 candidates challenged the long context comprehension of the model. A possible mitigation to this issue is replacing the $4k/2k$ selection step with Verbalized Maximum (Sec. 3), whose context only considers 2 documents at a time.
Table 7: Comparing logp over various retrieval models. In all cases, the embedding model that (pre-)selected kk/2k2k/$4k$ elements is Qwen3-embedding-8B. All other models are implied to be the number of parameters of Qwen3-family of models. We did not perform an exhaustive experiments over all parameters which leaves some cells empty, however our main claim holds.
hotpotQA CUAD $\phi=$ 0.6B $\phi=$ 32B $\phi=$ 0.6B $\phi=$ 32B $\psi$ $k=3$ 6 12 3 6 12 3 6 12 3 6 12 tfidf -10.6 -9.9 -9.1 -5.9 -5.4 -5.0 -309.4 -289.5 -269.0 -206.6 -191.3 -177.0 tfidf+MMR $\lambda=0.2$ -11.0 -11.0 -10.8 -6.2 -6.2 -6.2 -326.3 -325.4 -320.9 -219.3 -217.6 -211.6 tfidf+MMR $\lambda=0.5$ -10.6 -9.9 -9.4 -6.0 -5.7 -5.4 -319.2 -308.1 -299.9 -212.6 -202.8 -196.5 tfidf+MMR $\lambda=0.8$ -10.4 -9.7 -9.1 -5.8 -5.3 -5.0 -310.6 -289.6 -268.7 -206.9 -192.0 -177.1 bm25 -9.9 -9.7 -9.1 -5.4 -5.2 -5.0 -308.1 -288.3 -267.6 -205.6 -190.5 -175.7 bm25+MMR $\lambda=0.2$ -10.7 -10.6 -10.3 -6.0 -5.9 -5.7 -317.6 -315.6 -311.8 -212.0 -208.7 -204.4 bm25+MMR $\lambda=0.5$ -10.4 -10.1 -9.8 -5.9 -5.7 -5.5 -314.1 -307.1 -301.6 -208.7 -202.6 -197.4 bm25+MMR $\lambda=0.8$ -10.0 -9.6 -8.8 -5.6 -5.3 -4.9 -309.1 -295.3 -280.9 -205.9 -195.1 -185.1 embed(top-$k$) -9.6 -9.1 -8.6 -5.2 -5.1 -5.1 -291.0 -269.2 -253.9 -197.6 -180.9 -168.7 embed(top-kk)+MMR $\lambda=0.2$ -10.7 -10.6 -10.3 -5.7 -5.8 -5.7 -310.4 -305.9 -296.1 -208.9 -202.7 -193.8 embed(top-kk)+MMR $\lambda=0.5$ -9.7 -9.4 -8.6 -5.4 -5.4 -5.1 -305.8 -292.9 -279.1 -206.1 -194.6 -183.2 embed(top-kk)+MMR $\lambda=0.8$ -9.5 -9.2 -8.4 -5.2 -5.1 -5.0 -294.3 -276.1 -255.3 -199.2 -185.1 -169.2 embed(top-2k2k)+VA sort 0.6B -10.3 -9.9 -10.0 -301.5 -284.1 embed(top-4k4k)+VA sort 0.6B -11.0 -11.3 -10.7 -307.4 embed(top-2k2k)+VGSM 0.6B -9.6 -9.1 -8.6 -5.2 -5.1 -5.0 -290.5 -268.6 -251.3 -196.4 -180.4 -167.1 embed(top-4k4k)+VGSM 0.6B -9.5 -9.2 -10.0 -5.3 -5.2 -5.7 -289.8 -267.6 -250.3 -196.8 -179.6 -165.7 embed(top-2k2k)+VGSM 4B -9.2 -8.7 -8.2 -280.3 -267.3 -248.6 embed(top-4k4k)+VGSM 4B -9.0 -8.6 -8.2 -282.6 -265.0 -241.5 embed(top-2k2k)+VGSM 32B -9.2 -8.5 -7.9 -277.9 -261.7 -240.3 embed(top-4k4k)+VGSM 32B -8.6 -8.1 -7.8 -273.9 -250.4 -228.9
@@ 7 Related Work
Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023). Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary). Formalization approach to planning is also called a parse-and-solve model (Collins et al., 2022) as opposed to a LLM-as-Planner model that naively generates a plan with an LLM. To our knowledge, there is no work that simulates traditional graph search algorithms (e.g., A* (Hart et al., 1968), GBFS (Bonet & Geffner, 2001)) in a VA manner, i.e., by replacing all of its atomic operations, e.g., successor generation, heap-based priority queues, hash computation for duplicate detection, with LLM queries. Some work replaces some atomic operations through code generation: successor function and goal condition in Thought of Search (Katz et al., 2024) and collision checking in STPR (Djuhera et al., 2025).
A number of recent papers in the Algorithm with Predictions (Mitzenmacher & Vassilvitskii, 2022, ALPS) community achieved a significant success in augmenting algorithms by data-driven oracles. The particular strength of this line of work is the rigorous theoretical analysis of the error rate. VAs is a form of ALPS specialized in LLM-based oracles, thus combining ideas from ALPS community with VAs is a promising direction for future work.
Combining data-driven heuristics with existing algorithms in the classical AI tasks is also a common practice in the recent literature (e.g., classical planning (Núñez-Molina et al., 2024), boolean SAT (Duan et al., 2020), logistics tasks (Xin et al., 2021)).
8 Conclusion and Future Work
@@
While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc. Can Floyd’s tortoise-and-hare cycle detection algorithm for linked lists be verbalized to perform a tautology detection? Can garbage collection algorithms detect unimportant sentences in a paper draft? The potential of this paradigm is large and we invite various ideas for unique verbalizations and new applications.
@@ A.2 Kendall-Tau Score
The Kendall Tau correlation coefficient (Kendall, 1945) is a non-parameteric statistic that measures how much order is preserved between two sets of rankings. Kendall-Tau score has several ways to compute ($\tau_{a},\tau_{b},\tau_{c}$) depending on the existence of ties or the nature of the observations. Here we describe $\tau_{b}$ we used in our paper.
@@ B.1 Jaccard distance
B.2 LLM Query Prompts